OpenStudy (anonymous):

I am suppose to determine whether the statement below is true or false, and explain or support me answer by providing an example. For all real numbers a and b, 2a b = a2 + b2

OpenStudy (anonymous):

I need help bad

OpenStudy (anonymous):

\[2ab=a^2+b^2\]??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

almost anything would be a counter example you could try \(a=b=10\) and see that \[2\times 10\times 10\neq 10^2+10^2\] make up any numbers you choose, it is unlikely to work

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

yw